Astrid1:

Angles 1 and 2 are supplementary angles: Measure angle 1 = x-30 and measure of angle 2 = x+76, find the measure of angle 2 I'm pretty sure the first step is: x-30+x+76=180 But I'm stuck on step two.

Astrid1:

I believe step two is where you add "like terms" or whatever has the same variable. But there isn't a number before the variable. How would I go abt this?

Toaster:

Second step is to add all common factors on the left, then subtract or add common factors from the left to the right

Astrid1:

Yeah so I've gotten part of that done. So it's now 46+ somthing idk yet

Astrid1:

if there was another side of the equation this'd be easier for me. She never taught us equations with the number after the variable.

Toaster:

so you have 2x + 46=180 right?

jkmiec26:

Ok so I would say it would look something like this.  So you want to combine factors. Where you see I combined the x's and 76 and 30. Then you want to subtract 46 by 180 and divide and you get 67. You want to always check your answers at the end for better results.

Astrid1:

Yeah. My friends helping me rn lol

Astrid1:

Where'd you get 2x from though? is that jus m<2 as an unkown variable?

jkmiec26:

You want to combine the x's in the equation. x+x=2x.

jkmiec26:

Since you have 2 x's

Astrid1:

Oh okay.

Toaster:

x-30+x+76=180 one x before and one x after the -30

Astrid1:

Thanks for your help guys! I'm keep this in my notes for extra reference so this doesn't happen again. 🤭

jkmiec26:

lmaoo ofc